OpenStudy (diyadiya):

If α≠β but \(α^2=5α−3\) ,\(β^2=5β−3\), then the equation whose roots are \(\frac{\alpha}{\beta}\) and \(\frac{\beta}{\alpha}\)

OpenStudy (diyadiya):

I have to find the equation

OpenStudy (ash2326):

@Diyadiya did you find \(\alpha\) and \(\beta\) ??

OpenStudy (anonymous):

You have to really? really? what if you won't ?

OpenStudy (lgbasallote):

ffm to the rescue!!! \m/ hahaha

OpenStudy (diyadiya):

@ash2326 Nope i didnt

OpenStudy (diyadiya):

\[\alpha+\beta=5\]

OpenStudy (ash2326):

\(\alpha \) \(\ne\) \(\beta\) There are two roots of \[\alpha^2=5\alpha-3\] one roots is \(\alpha\) and another is \(\beta\) so \[\alpha+\beta=5\] and \[\alpha\beta=3\]

OpenStudy (anonymous):

Your required equation is \(3x^2 -19x+3=0 \)

OpenStudy (diyadiya):

Sorry i was afk

OpenStudy (ash2326):

so the equation whose roots are \(\frac{\alpha}{\beta}\) and \(\frac{\beta}{\alpha}\) is \[ x^2-(\frac{\alpha}{\beta}\ + \frac{\beta}{\alpha})+\frac{\alpha}{\beta} \times \frac{\beta}{\alpha}=0\]

OpenStudy (diyadiya):

@FoolForMath yeah!

OpenStudy (diyadiya):

okay Thank you :) @ash2326

OpenStudy (ash2326):

Welcome:D

OpenStudy (anonymous):

Do you read Mathematics today?

OpenStudy (diyadiya):

No ?

OpenStudy (diyadiya):

Why?

OpenStudy (anonymous):

You should, lots of cute problems in there.

OpenStudy (diyadiya):

lol ok

OpenStudy (diyadiya):

I still didn't get it how do you find the sum in \[x^2-(\frac{\alpha}{\beta}\ + \frac{\beta}{\alpha})+\frac{\alpha}{\beta} \times \frac{\beta}{\alpha}=0 \]

OpenStudy (anonymous):

Do you know: \(a^2+b^2 = (a+b)^2-2 ab \)?

OpenStudy (diyadiya):

Yeah

OpenStudy (anonymous):

then use it ;)

OpenStudy (diyadiya):

Ok

OpenStudy (diyadiya):

Thanks

OpenStudy (anonymous):

Welcome :) Btw, Sometimes it's good to use that cute head of yours :P I mean Leena's :D